# Enhancing variational quantum algorithms by balancing training on classical and quantum hardware

### v1

@@ Appendix B Example instances of $\mathfrak{g}$-sim reaching accurate solution without quantum hardware
Figure 11: Error vs Iteration curve for finding ground state of an XY Hamiltonian using only $\mathfrak{g}$-sim . For 13 (11(a)) and 17 (11(b)) qubits, $\mathfrak{g}$-sim is enough to obtain the ground state and does not require our protocol.
Appendix C VQE for TFIM Hamiltonian
@@
$H_{TFIM}=\sum_{i}\alpha_{i}X_{i}X_{i+1}+\sum_{j}\beta_{j}Z_{j}$ (23)
which has a poly-DLA scaling as $2n^{2}-n$ for $n$ qubits. We measure the improvement in terms of Error, Success and QPU calls reduction as defined in IV.
We tabulate the numerical results comparing $\mathfrak{g}$-sim , Full-PSR, Alternate and Alternate+Simultaneous in Table 4 and 5 and provide plots for visualization in Fig.12. Except the 12 qubit case, where $\mathfrak{g}$-sim is able to produce accurate solutions (similar to the scenario mentioned in App.B), our proposed Alternate+Simultaneous method generally achieves a lower error and higher success rate compared to the other protocols, while almost always reducing the QPU calls necessary compared to Full-PSR.
Figure 12: TFIM Hamiltonian VQE: The success (Column 1), error (Column 2) and QPU calls reduction (Column 3) are plotted for configurations with 1, 3, 6, and 9 layers of YZ linear ansatz in Block Q and Hamiltonian DLA gates in Block G, for various qubit counts in each row. For each configuration and metric, PSR (orange), Alternate(blue) and Alternate+Simultaneous(purple) training methods are compared. For error (Column 2), we plot the median and the $25\%$ and $75\%$ quantiles (in black). For QPU calls reduction only Alternate and Alternate+Simultaneous is plotted since we compare it relative to PSR (which will coincide with 0% QPU calls reduction). The spread in standard deviation is shown in black. We consistently see reduction in QPU calls using our methods, with frequently lower error values and higher success metrics as compared to PSR
Figure 13: Gradients for TFIM Hamiltonian VQE: Variance of the first parameter from PSR and $\mathfrak{g}$-sim block are plotted with increasing qubits for both Full-PSR (blue) and our Alternate method (orange) for finding ground state of TFIM Hamiltonian using VQE. The diagrams also show the plots for varying number of PSR block layers (1,3,6,9 from top to bottom). For comparison, the same task is performed using a deep HEA (referred as HEA50) with standard PSR and variance of its first partial derivative is plotted in green. The green curve is the same in all of the plots and provided as a reference for our improved gradients. The plot shows that our chosen ansatz is able to preserve higher magnitude of gradients even at high qubits in both of the blocks.
@@
Table 4: TFIM Hamiltonian (Success and QPU Reduction): The success and QPU calls reduction are shown for the configurations with 1, 3, 6, and 9 layers of YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G using Full-PSR, Alternate and Alternate + Simultaneous training (referred as Alt+Sim). The QPU call reduction is measured relative to the Full-PSR method where all parameters are trained by standard PSR. The metrics show that our methods are generally better at reaching higher success rates and lower error(in Table 5), while the QPU calls are reduced significantly
Qubits Error (1e-5)
@@
(a) Error for 1 layer YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G
Qubits Error (1e-5)
@@
(b) Error for 3 layer YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G
Qubits Error (1e-5)
@@
(c) Error for 6 layer YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G
Qubits Error (1e-5)
@@
(d) Error for 9 layer YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G
@@
Table 5: TFIM Hamiltonian (Error): The error median, 25% quantile (Q25) and 75% quantile (Q75) for Full-PSR, Alternate and Alternate + Simultaneous (referred as Alt+Sim) for the configurations with 1, 3, 6, and 9 layers of YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G. The error for $\mathfrak{g}$-sim with only TFIM Hamiltonian DLA gates for each qubit count is shown for comparison, which is several orders of magnitude higher than the other methods. Alternate and Alternate + Simultaneous method shows generally lower errors compared to the Full-PSR training
@@
We also provide plots of variance of gradients for the first parameter from each of $\mathfrak{g}$-sim and PSR block in Fig. 13. Consistent with our previous example, the gradients decay slower using our choice of ansatz as compared to a deep Hardware Efficient Ansatz with 50 layers.

### v2

@@ Appendix G Example instances of $\mathfrak{g}$-sim reaching accurate solution without quantum hardware
Figure 13: Relative Error vs Iteration curve for finding ground state of an XY Hamiltonian using only $\mathfrak{g}$-sim . For 13 (13(a)) and 17 (13(b)) qubits, $\mathfrak{g}$-sim is enough to obtain the ground state and does not require our protocol.
Appendix H VQE for TFIM Hamiltonian
@@
$H_{TFIM}=\sum_{i=0}^{N}\alpha_{i}X_{i}X_{i+1}+\sum_{j}\beta_{j}Z_{j}$ (26)
which has a poly-DLA scaling as $2n^{2}-n$ for $n$ qubits. We measure the improvement in terms of Relative Error, Success and QPU calls reduction as defined in IV.
We tabulate the numerical results comparing $\mathfrak{g}$-sim , Full-PSR, Alternate and Alternate+Simultaneous in Table 5 and 6 and provide plots for visualization in Fig. 17. Except the 12 qubit case, where $\mathfrak{g}$-sim is able to produce accurate solutions (similar to the scenario mentioned in App.G), our proposed Alternate+Simultaneous method generally achieves a lower relative error and higher success rate compared to the other protocols, while almost always reducing the QPU calls necessary compared to Full-PSR.
(a) Success, Relative Error and QPU Call Reduction for 1 YZ linear layers in Block Q
(a) Success, Relative Error and QPU Call Reduction for 3 YZ linear layers in Block Q
(a) Success, Relative Error and QPU Call Reduction for 6 YZ linear layers in Block Q
(a) Success, Relative Error and QPU Call Reduction for 9 YZ linear layers in Block Q
Figure 17: TFIM Hamiltonian VQE (1,3,6 and 9 YZ linear layer): The success (Row 1), relative error (Row 2) and QPU calls reduction (Row 3) are plotted for configurations with 1 (a), 3 (b) and 6 (c) and 9 (d) layers of YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G for 8 to 14 qubits. For each qubit count and metric, PSR (orange), Alternate (blue) and Alternate+Simultaneous(purple) training methods are compared. On the first row, we compare the success rates for both Alternate and Alternate+Simultaneous protocol with PSR. For relative error (Row 2), we plot the median and the $25\%$ and $75\%$ quantiles (in black). The Alt+Sim method gives comparable or lower relative error than PSR for all qubits. Most importantly, on the lowermost row QPU calls reduction is plotted for only Alternate and Alternate+Simultaneous since we compare it relative to PSR (which will coincide with 0% QPU calls reduction). The spread in standard deviation is shown in black. Except for 14 qubit case in 9 layer configuration, we consistently see reduction in QPU calls using our methods, with lower relative error values and higher success metrics as compared to PSR.
(a) 1 YZ linear layer in Block Q
(b) 3 YZ linear layer in Block Q
(c) 6 YZ linear layer in Block Q
(d) 9 YZ linear layer in Block Q
Figure 18: Gradients for TFIM Hamiltonian VQE: Variance of the first parameter from Block Q ($U_{q}$,red) and G ($U_{g}$,blue) are plotted with increasing qubits for both Full-PSR and our Alternate method for finding ground state of TFIM Hamiltonian using VQE. The diagrams also show the plots for 1 (a), 3 (b), 6(c), 9 (d) YZ linear layers in Block Q . For comparison, the same task is performed using a deep HEA (referred as HEA50) with standard PSR and variance of its first partial derivative is plotted in green. The green curve is the same in all of the plots and provided as a reference for our improved gradients. The plot shows that our chosen ansatz is able to preserve higher magnitude of gradients even at high qubits in both of the blocks.
@@
Table 5: TFIM Hamiltonian (Success and QPU Reduction): The success and QPU calls reduction are shown for the configurations with 1, 3, 6, and 9 layers of YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G using Full-PSR, Alternate and Alternate + Simultaneous training (referred as Alt+Sim). The QPU call reduction is measured relative to the Full-PSR method where all parameters are trained by standard PSR. The metrics show that our methods are generally better at reaching higher success rates and lower relative error(in Table 6), while the QPU calls are reduced significantly
Qubits Relative Error (1e-5)
@@
(a) Relative Error for 1 layer YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G
Qubits Relative Error (1e-5)
@@
(b) Relative Error for 3 layer YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G
Qubits Relative Error (1e-5)
@@
(c) Relative Error for 6 layer YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G
Qubits Relative Error (1e-5)
@@
(d) Relative Error for 9 layer YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G
@@
Table 6: TFIM Hamiltonian (Relative Error): The relative error median, 25% quantile (Q25) and 75% quantile (Q75) for Full-PSR, Alternate and Alternate + Simultaneous (referred as Alt+Sim) for the configurations with 1, 3, 6, and 9 layers of YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G. The relative error for $\mathfrak{g}$-sim with only TFIM Hamiltonian DLA gates for each qubit count is shown for comparison, which is several orders of magnitude higher than the other methods. Alternate and Alternate + Simultaneous method shows generally lower error compared to the Full-PSR training
@@
We also provide plots of variance of gradients for the first parameter from each of Block Q and G in Fig. 18. Consistent with our previous example, the gradients decay slower using our choice of ansatz as compared to a deep Hardware Efficient Ansatz with 50 layers.
Output: $\theta^{*},\phi^{*}$
Initialize: $\vec{\theta}_{0}\sim\mathcal{N}(0,1);i=0$
begin
       poly-DLA Phase
3       Randomly initialize $\vec{\phi}_{0}\sim\mathcal{N}(0,1)$
4       Define the cost function: $C_{poly}(\vec{\theta},\vec{\phi})=\langle 0^{\otimes n}|U_{q}^{\dagger}(\vec{\theta})U_{g}^{\dagger}(\vec{\phi})H_{poly}U_{\mathfrak{g}}(\vec{\phi})U_{q}(\vec{\theta})|0^{\otimes n}\rangle$
5       Run Alternate optimization given in Alg.1 for $T_{alt}$ iterations on cost function $C_{poly}(\vec{\theta},\vec{\phi})$
6       Run Simultaneous optimization given in Alg.2 for $T_{sim}$ iterations on cost function $C_{poly}(\vec{\theta},\vec{\phi})$
7       Extract the optimized parameters $(\vec{\theta}^{*},\vec{\phi}^{*})$
       Exponential-DLA Phase
8       Define the full cost function, $C_{full}(\vec{\theta},\vec{\phi})=\langle 0^{\otimes n}|U_{q}^{\dagger}(\vec{\theta})U_{g}^{\dagger}(\vec{\phi})HU_{\mathfrak{g}}(\vec{\phi})U_{q}(\vec{\theta})|0^{\otimes n}\rangle$ initialized at $(\vec{\theta}^{*},\vec{\phi}^{*})$
12
Algorithm 4 VQE for General Hamiltonian
Appendix I VQE for LTFIM Hamiltonian (exponential-DLA)
In this section, we provide the algorithm (Algo. 4) and plots for VQE on 10 and 12 qubit LTFIM Hamiltonian using PSR and our proposed method in Fig. 19. The relative error metric in Eq.(7) is shown for 3 and 6 YZ linear layers in Block Q. Although the initial training phase is based on the reduced Hamiltonian, the plotted energy values correspond to expectation values of the full Hamiltonian.
Experiment Details: $U_{q}$ is composed of 3 layers of YZ linear ansatz, while $U_{\mathfrak{g}}$ is generated out of gates from $\mathfrak{g}_{TFIM}$. In the shown examples, we run Alternate optimization for 250 iterations and Simultaneous optimization for 100 iterations with the reduced Hamiltonian, followed by 200 iterations of $U_{q}$ and 1000 iterations of full circuit training. Similarly for Full-PSR training, we run 1450 iterations of PSR to match the number of total iterations in both methods. However, these choices are made ad-hoc and can be modified based on other stopping conditions, such as early stopping to further reduce QPU calls.
(a) Trained using standard PSR
(b) Pre-training using our method followed by PSR
(c) Trained using standard PSR
(d) Pre-training using our method followed by PSR
(e) Trained using standard PSR
(f) Pre-training using our method followed by PSR
Figure 19: Demonstration of PSR (a),(c), (e) and our method (b),(d),(f) to run VQE for 10 and 12 qubit LTFIM Hamiltonian (with an exponential DLA) with 3 and 6 YZ linear layers in Block Q. The orange and green regions show Alternate and Simultaneous training respectively for the restricted Hamiltonian, followed by training Block Q and full circuit in blue and pink regions respectively for the full LTFIM Hamiltonian. We notice that our training methods in the initial phase allows VQE to reach much closer to the target ground state during the final training, as compared to just using PSR for the full training, for similar number of of total training iterations. For the overall training, we also reduce the QPU calls ($\sim 6-10\%$).
As shown in Fig.19, the algorithm is indeed able to converge to good solutions by using the our protocols in the first phase of the training. Compared to PSR, our protocol shows lower relative error value on an average. The QPU calls required for the 10-qubit Hamiltonian VQE using PSR is $54.5e+$ 5 and $56.3e+$ 5 for Layer 3 and 6 respectively, while our proposed method requires $54.065e+$ 5 ($\sim 9.6\%$ reduction) and $55.925e+$ 5 ($\sim 6.5\%$ reduction) QPU calls respectively. Similarly, for 12 qubit, PSR uses $58.28e+$ 5 for Layer 6 while our method requires $57.698e+$ 5 ($\sim 7\%$ reduction) respectively.
Appendix J Experimental details of VQE on LiH and Classification task
LiH Experimental details: $U_{q}$ is composed of 6 layers of YZ linear ansatz, while $U_{\mathfrak{g}}$ is generated out of gates from the above DLA. In the shown examples, we run Alternate optimization for 250 iterations and Simultaneous optimization for 100 iterations with the reduced Hamiltonian, followed by 200 iterations of $U_{q}$ and 1000 iterations of full circuit training. Similarly for Full-PSR training, we run 1450 iterations of PSR to match the number of total iterations in both methods. These ad-hoc choices can potentially be further optimized to reduce resources and improve accuracy.
Classification Experiment Details: We construct HELIA with 9 YZ linear layers in $U_{q}$. $U_{\mathfrak{g}}$ is constructed from one of the 3 DLA choices mentioned above, and we repeat the experiment for each of the choices. The measurement operator is chosen randomly from the corresponding DLA and fixed at the beginning of the experiment. For each of these configurations of DLA, we run 20 independent trials and compare test accuracy.